After five science tests, Sonya's average score is 89. If she gets a 95 on the sixth test, what is her new average? I need some explanation on how to get this answer.
Answer by fcabanski(1391)

The formula for average is sum / (number of numbers) = avg.


Sum / 5 = 89


sum (of the five numbers) = 89*5 = 445


After the 6th test, a 95, the new sum is 445 + 95 = 540. The new average is 540/6 = 90.